# SHIELD: LLM-Driven Schema Induction for Predictive Analytics in EV Battery Supply Chain Disruptions

### v1

@@ Abstract
Zhi-Qi Cheng1  Yifei Dong2  Aike Shi3  Wei Liu4  Yuzhi Hu5 Jason O’Connor1Alexander Hauptmann1Kate Whitefoot1 1Carnegie Mellon University  2Columbia University  3Georgia Institute of Technology 4University of Michigan, Ann Arbor  5Boston University Project Page: https://f1y1113.github.io/MFI/
{NoHyper}Completed by Y. Dong and Y. Hu during their remote visit, and by A. Shi and W. Liu during their internships at CMU. Z. Cheng, Y. Dong, A. Shi, W. Liu, and Y. Hu contributed equally. J. O’Connor, A. Hauptmann, and K. Whitefoot provided guidance. Detailed contributions are in Appx. L.
@@ 1 Introduction
2 Schema Learning for EV Supply Chains
@@
3 Dynamic Disruption Analysis
@@
Algorithm 1 Supply Chain Disruption Prediction
2:Output: Refined predictions $\hat{y}^{\prime}$
4:for $l=1$ to $L$ do
6:end for
7:$\hat{y}\leftarrow\mathbf{H}^{(L)}$
9:for each prediction $\hat{y}_{i}$ do
10:     $\hat{y}_{i}^{\prime}\leftarrow\text{Constrain}(\hat{y}_{i})$
12:end for
13:Coreference Resolution $\triangleright$ Link related events
14:for each pair of events $(E_{i},E_{j})$ do
16:     if $R_{ij}$ is coreferential then
17:         Link $E_{i}$ and $E_{j}$
18:     end if
19:end for
20:Return: Refined predictions $\hat{y}$
@@
4 Experiments
4.1 Schema Learning Performance
4.2 Disruption Detection Performance
@@ 5 Conclusion
Appendix A Related Work
@@ I.1 Experiment Setup
In Experiment 4.1, we evaluate the efficacy of three distinct Large Language Models (LLMs) in extracting hierarchical structures from our schema learning dataset. Leveraging domain expert knowledge, we annotate individual schemas for each article in our academic corpus using our proprietary system viewer and editor. We then employ the methodology outlined in Appx. D to synthesize these schemas into an integrated library. This combination of individual schemas and the integrated library serves as the ground truth for our hierarchical information extraction phase. Our schema learning performance evaluation consists of two key components. First, we compare the hierarchical information extracted by the three LLMs against our established ground truth. Second, we assess the consistency, accuracy, and completeness of the hierarchical structures derived from the textual content of each article in the schema learning dataset, with domain experts actively participating in this evaluation process.
In Experiment 4.2, we apply a similar annotation methodology to our news dataset as used for the schema learning dataset. However, annotating the news dataset presents unique challenges, as news reports typically do not explicitly elucidate the connections between events. Instead, they often employ speculative language to describe event interrelations. To ensure annotation accuracy, we heavily rely on domain knowledge derived from scenario documents throughout the annotation process. Subsequently, we utilize the ground truth extracted from these reports to evaluate our system’s performance in predicting news report outcomes.
@@ I.2 Evaluation Metrics
J.1 Case 1: Lithium Supply Chain Disruption in Early 2023
@@
J.2 Case 2: Nickel and Cobalt Supply Issues in March 2023
@@
J.3 Case 3: Impact of the Inflation Reduction Act in August 2023
In August 2023, the United States passed the Inflation Reduction Act, which included significant incentives for domestic EV battery production. This led to a rapid increase in investments but also highlighted potential material shortages, causing disruptions in the EV battery supply chain.
@@
The SHIELD user interface is designed to be intuitive and user-friendly, facilitating the efficient upload and analysis of news reports. It comprises two main sections: the news report upload and the disruption analysis results.
@@
  • Yifei Dong: Designed and modified prompts for generating structured information, designed the format for structured information, wrote scripts for converting structured information to SDF, and for generating structured information using Llama3 (partially contributed by Aike Shi).
  • Yifei Dong, Aike Shi, and Wei Liu: Generated structured information using GPT-4o with zero-shot learning.
@@
  • Yifei Dong: Performed system implementation and debugging (with minor contributions from Aike Shi).
@@
  • Yifei Dong: Provided the design.
  • Aike Shi and Wei Liu: Implemented interface.
  • Wei Liu: Optimized the interface.
@@
  • Yifei Dong: Contributed to all sections except the news dataset part.
@@
  • Zhi-Qi Cheng: Organized and rewrote the paper, and supervised the modification of all figures.
@@
  • Kate Whitefoot and Alexander Hauptmann: Provide guidance and supervision for the entire project.

### v2

@@ Abstract
Zhi-Qi Cheng1  Yifei Dong2  Aike Shi3  Wei Liu4  Yuzhi Hu5 Jason O’Connor1Alexander G. Hauptmann1Kate S. Whitefoot1 1Carnegie Mellon University  2Columbia University  3Georgia Institute of Technology 4University of Michigan, Ann Arbor  5Boston University Project Page: https://f1y1113.github.io/MFI/
{NoHyper}Completed by Y. Dong and Y. Hu during remote visits, and A. Shi and W. Liu during CMU internships. Z. Cheng, Y. Dong, A. Shi, W. Liu, and Y. Hu contributed equally. J. O’Connor, A. Hauptmann, and K. Whitefoot provided guidance. See Appx. L for details. Correspondence: zhiqic,alex@cs.cmu.edu, kwhitefoot@andrew.cmu.edu.
@@ 1 Introduction
2 Related Work
Supply Chain Risk Management. AI has been increasingly applied to predict and manage supply chain risks Ganesh and Kalpana (2022). Agent-based approaches Pino et al. (2010); Giannakis and Louis (2011, 2016); Blos et al. (2015) facilitate inter-agent communication for forecasting but often suffer from limited predictive power and parameter constraints. Rule-based methods Gallab et al. (2019); Behret et al. (2012); Paul (2015); Paul et al. (2017); Awasthi et al. (2018); Camarillo et al. (2018) offer decision frameworks with minimal quantitative insights. Machine Learning (ML) and Deep Learning (DL) techniques have improved forecasting and disruption prediction Silva et al. (2017); Hegde and Rokseth (2020); Garvey et al. (2015); Ruz et al. (2020); Aljohani (2023); Carbonneau et al. (2008), yet many focus on predictive performance at the expense of interpretability Hendriksen (2023); Makridis et al. (2023). Recent work has explored large language models (LLMs) in supply chain management Ray (2023), but interpretability remains a challenge. Our approach integrates LLMs to enhance both predictive accuracy and interpretability by extracting hierarchical knowledge-graph structures to forecast disruptions.
Schema Induction & Learning. Building on early schema induction work Anderson et al. (1979); Evans (1967), LLMs Brown et al. (2020); Rae et al. (2021) have shown strong schema-learning abilities with minimal supervision. Recent strategies, such as contextual explanations Wei et al. (2021); Lampinen et al. (2022), rationale-augmented models Wang et al. (2022b), and incremental prompting Li et al. (2023), have further refined schema induction. Transformer-based methods Li et al. (2020, 2021) excel at schema representation through graph structures, with human feedback playing a vital role in improving model accuracy Mondal et al. (2023); Yang et al. (2024); Zhang et al. (2023). Our method leverages these advancements, combining human feedback with LLM-driven schema induction to improve accuracy and relevance in disruption prediction.
Event Extraction & Analysis. Event extraction has evolved from handcrafted features Ahn (2006) to neural models like recurrent Nguyen et al. (2016); Sha et al. (2018), convolutional Chen et al. (2015), graph Zhang and Ji (2021), and transformer-based networks Liu et al. (2020). Advances in argument extraction Wang et al. (2019), zero-shot learning Huang et al. (2018), and weak supervision Chen et al. (2015) have boosted performance. Our approach enhances event extraction by using fine-tuned RoBERTa models and graph convolutional networks (GCNs) to capture complex event relationships and cascading effects, offering deeper insights into supply chain disruptions compared to traditional methods.
3 Schema Learning for Supply Chain Disruptions
@@
4 Dynamic Analysis of Supply Chain Disruptions
@@
Algorithm 1 Supply Chain Disruption Prediction
2:Output: Refined predictions $\hat{y}^{\prime}$
4:for $l=1$ to $L$ do
6:end for
7:$\hat{y}\leftarrow\mathbf{H}^{(L)}$
9:for each prediction $\hat{y}_{i}$ do
10:     $\hat{y}_{i}^{\prime}\leftarrow\text{Constrain}(\hat{y}_{i})$
12:end for
13:Coreference Resolution $\triangleright$ Link related events
14:for each pair of events $(E_{i},E_{j})$ do
16:     if $R_{ij}$ is coreferential then
17:         Link $E_{i}$ and $E_{j}$
18:     end if
19:end for
20:Return: Refined predictions $\hat{y}$
@@
5 Experiments
5.1 Schema Learning Performance
5.2 Disruption Detection Performance
@@ 5.3 Qualitative Analysis & User Interface
5.4 Disruption Prediction Case Studies
Our system effectively forecasted key supply chain disruptions, providing insights that enabled stakeholders to take proactive actions. For example, following the passage of the Inflation Reduction Act (2022), which incentivized domestic EV battery production, our system predicted potential material shortages. By analyzing shifts in global material flows and the effects of policy changes on supply-demand dynamics, it enabled early interventions to minimize risks. Similarly, during geopolitical tensions between Australia and China in 2023, our system identified vulnerabilities in the lithium supply chain by monitoring export data and geopolitical developments, helping stakeholders adapt their sourcing strategies in time. In another instance, the system anticipated cobalt supply issues resulting from labor strikes and regulatory changes in the Democratic Republic of Congo (2023), allowing companies to diversify sources and increase inventory buffers. These cases, detailed further in Appendix J, illustrate how data-driven predictions enhance supply chain resilience and support timely decision-making in a volatile global market.
@@ 6 Conclusion
Acknowledgments
This research was supported by the Carnegie Mellon Manufacturing Futures Institute and the Manufacturing PA Innovation Program. The authors also thank the School of Computer Science (SCS) at Carnegie Mellon University, particularly the High Performance Computing (HPC), for providing essential computational resources. The views expressed are those of the authors and do not necessarily reflect those of the funding agencies.
Appendix A Extended Related Work
@@ I.1 Experiment Setup
In Experiment 5.1, we evaluate the efficacy of three distinct Large Language Models (LLMs) in extracting hierarchical structures from our schema learning dataset. Leveraging domain expert knowledge, we annotate individual schemas for each article in our academic corpus using our proprietary system viewer and editor. We then employ the methodology outlined in Appx. D to synthesize these schemas into an integrated library. This combination of individual schemas and the integrated library serves as the ground truth for our hierarchical information extraction phase. Our schema learning performance evaluation consists of two key components. First, we compare the hierarchical information extracted by the three LLMs against our established ground truth. Second, we assess the consistency, accuracy, and completeness of the hierarchical structures derived from the textual content of each article in the schema learning dataset, with domain experts actively participating in this evaluation process.
In Experiment 5.2, we apply a similar annotation methodology to our news dataset as used for the schema learning dataset. However, annotating the news dataset presents unique challenges, as news reports typically do not explicitly elucidate the connections between events. Instead, they often employ speculative language to describe event interrelations. To ensure annotation accuracy, we heavily rely on domain knowledge derived from scenario documents throughout the annotation process. Subsequently, we utilize the ground truth extracted from these reports to evaluate our system’s performance in predicting news report outcomes.
@@ I.2 Evaluation Metrics
J.1 Case 1: Impact of the Inflation Reduction Act in August 2022
In August 2022, the United States passed the Inflation Reduction Act, which included significant incentives for domestic EV battery production. This led to a rapid increase in investments but also highlighted potential material shortages, causing disruptions in the EV battery supply chain.
J.2 Case 2: Lithium Supply Chain Disruption in Early 2023
@@
J.3 Case 3: Nickel and Cobalt Supply Issues in March 2023
@@
The SHIELD user interface is designed to be intuitive and user-friendly, facilitating the efficient upload and analysis of news reports.
@@
  • Yifei Dong and Aike Shi: Designed and modified prompts for generating structured information, designed the format for structured information, wrote scripts for converting structured information to SDF, and for generating structured information using Llama3.
  • Yifei Dong, Aike Shi, Wei Liu, and Yuzhi Hu: Generated structured information using GPT-4o with zero-shot learning.
@@
  • Yifei Dong and Aike Shi: Performed system debugging and testing.
@@
  • Aike Shi and Yifei Dong: Provided the design.
  • Wei Liu: Optimized the design.
  • Aike Shi: Implemented the interface.
@@
  • Yifei Dong: Contributed to schema learning parts.
@@
  • Zhi-Qi Cheng: Organized and rewrote the entire paper, and supervised the modification of all figures.
Presentation and Project Webpage:
  • Aike Shi and Zhi-Qi Cheng: Presented the paper.
  • Yifei Dong: Built the project website.
  • Aike Shi: Created the video demo.
  • Wei Liu and Yuzhi Hu: Provided materials for the project website.
@@
  • Zhi-Qi Cheng, Kate S. Whitefoot, and Alexander G. Hauptmann: Provided guidance and supervision for the entire project.